3. Given: ΝΣ, Χ, y – Σ., & ΣΝ1y Β = Σ Α Σ., και - Σε, 8, Σ, 8, 9 Q = NΣ, x? - (Σ, x) Show all calculations using the above

How would I use these formulas to find the answer (answer is [3.075]).
Answer #1

Ans Q=NÉX-C xoa No 4 (number of points.) Q = 4(x,+ + x + x3 + xy?) - (x,+ x2 + x3+ xy) ? Q=4(1.32+2.52 +4.92+6012) - (1-3 +2.